prove that tanA + 2tan2A + 4tan4A + 8cot8A = cotA

prove that


tanA + 2tan2A + 4tan4A + 8cot8A = cotA

Grade:11

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Paradox

Solving from right
tan A + 2tan2A + 4tan4A + 8cot8A = tan A + 2tan2A + 4tan4A + 8/tan8A
= tan A + 2tan2A + 4tan4A + 8(1-tan24A)/2tan4A
= tan A + 2tan2A + [{4tan4A(tan4A)} + 4 (1-tan24A)]/tan4A
= tan A + 2tan2A + [4tan24A + 4 - tan24A]/tan4A  
= tan A + 2tan2A + 4/tan4A

Proceeding as above, we will reach to
tan A + 2tan2A + 4tan4A + 8cot8A = 1/tanA = cotA
Hence proved.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

TANA + 2TAN2A + 4TAN4A + 8COT8A = COTA .........1

WE HAVE FORMULA , COTA-TANA=2COT2A

 

EQ 1 CAN BE WRITTEN AS

   2TAN2A + 4TAN4A + 8COT8A = COTA-TANA

  2TAN2A + 4TAN4A + 8COT8A = 2COT2A               ( BY USING FORMULA )

  4TAN4A + 8COT8A = 2COT2A - 2TAN2A              

   4TAN4A + 8COT8A =4COT4A                               (BY USING FORMULA)

               8COT8A = 4COT4A - 4TAN4A

                            =8COT8A                        (BY USING FORMULA )

HENCE PROVED

 

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Solving from right
tan A + 2tan2A + 4tan4A + 8cot8A
= tan A + 2tan2A + 4tan4A+ 8/tan8A
= tan A + 2tan2A + 4tan4A + 8(1-tan24A)/2tan4A
= tan A + 2tan2A + [{4tan4A(tan4A)} + 4 (1-tan24A)]/tan4A
= tan A + 2tan2A + [4tan24A + 4 - tan24A]/tan4A
= tan A + 2tan2A + 4/tan4A
Proceeding as above, we will reach to
tan A + 2tan2A + 4tan4A + 8cot8A = 1/tanA = cotA

Thanks and Regards